# GRAPES: Learning to Sample Graphs for Scalable Graph Neural Networks

### v2

@@ Abstract
Graph neural networks (GNNs) learn to represent nodes by aggregating information from their neighbors. As GNNs increase in depth, their receptive field grows exponentially, leading to high memory costs. Several existing methods address this by sampling a small subset of nodes, scaling GNNs to much larger graphs. These methods are primarily evaluated on homophilous graphs, where neighboring nodes often share the same label. However, most of these methods rely on static heuristics that may not generalize across different graphs or tasks. We argue that the sampling method should be adaptive, adjusting to the complex structural properties of each graph. To this end, we introduce GRAPES, an adaptive sampling method that learns to identify the set of nodes crucial for training a GNN. GRAPES trains a second GNN to predict node sampling probabilities by optimizing the downstream task objective. We evaluate GRAPES on various node classification benchmarks, involving homophilous as well as heterophilous graphs. We demonstrate GRAPES’ effectiveness in accuracy and scalability, particularly in multi-label heterophilous graphs. Unlike other sampling methods, GRAPES maintains high accuracy even with smaller sample sizes and, therefore, can scale to massive graphs. Our code is publicly available at https://github.com/dfdazac/grapes.
@@ 1 Introduction
In many applications, data is represented with graph structures, such as in recommender systems, social networks, and the chemical and medical domains [31, 42, 25]. In these domains, graph neural networks (GNNs) are a powerful tool for representation learning on graphs [23, 38, 46].
Despite the broad range of applications of GNNs, scalability remains a significant challenge [35]. Unlike traditional machine learning problems where data is assumed to be i.i.d., the graph structure introduces dependencies between a node and its neighborhood. This complicates partitioning data into mini-batches. Additionally, the number of nodes that a GNN needs to process increases exponentially with the number of layers.
Most current graph sampling methods use a fixed heuristic to compute node inclusion probabilities: The sampling process is independent of the node features, graph structure, or the GNN task performance [9, 51, 47]. For simple graphs like homophilous graphs [50, 49, 48] where there is a strong correlation between labels of a node and its neighbors, a fixed heuristic typically suffices. However, these methods fall short when applied to graphs in which only a few of the nodes in the neighborhood provide relevant information for the task at hand. Consequently, we study adaptive sampling, where the sampling method adapts dynamically to the task by learning which nodes should be included. Our adaptive sampling technique learns to sample by directly minimizing the classification loss of the downstream GNN. This differs from most existing graph sampling methods, which aim to approximate the full-batch GNN behavior [9, 51, 47, 18], opting for an indirect approach to achieve high accuracy.
@@
To enable backpropagation through the sampling step, we need a gradient estimator. We compare a reinforcement learning (RL) approach and a GFlowNet (GFN) approach [6]. Both approaches allow us to consider the sampling process and the GNN computation together and to train them concurrently. This allows the sampler to adapt based on factors like the node features, the graph structure, the sample size, and other contextual features.
In addition to the datasets commonly used in the literature, we evaluate GRAPES on several heterophilous and/or multi-label graphs and demonstrate its effectiveness on these complicated graphs. Moreover, we spend significant effort on the evaluation protocol for fairly testing graph sampling methods. In particular, we ensure that all sampling methods are evaluated under the same conditions and on the same GNN architecture to eliminate any confounding factors and to ensure that any changes in performance can only be attributed to the sampling method. To the best of our knowledge, we are the first to perform such a rigorous comparison on twelve varied datasets.
@@
    Competitive Performance. GRAPES is competitive to other sampling methods in accuracy across datasets. On heterophilous and multi-label datasets GRAPES provides state-of-the-art performance.
@@
    Homophilous Graphs. For some benchmarks, mostly those based on homophilous graphs, simple baselines like random sampling perform as well as any state-of-the-art method, including GRAPES.
@@
    Memory Efficiency.  GRAPES is competitive with a state-of-the-art non-sampling GNN scaling method, while using up to an order of magnitude less GPU memory.
@@ 2 Related Work
Fixed Sampling Policy: In this category, the sampling policy is independent of the training of the GNN and is based on a fixed heuristic that does not involve any training. Given a set of target nodes, i.e. the nodes that are to be classified, node-wise sampling methods sample a given number of nodes for each target node. GraphSage [16] is a node-wise sampling method that randomly samples nodes. However, node-wise sampling can result in nodes being sampled multiple times redundantly because it can be the neighbor of several nodes [51]. A more efficient approach is layer-wise sampling, for example, FastGCN [9] and LADIES [51]. They aim to minimize variance by sampling nodes in each layer with probabilities proportional to their degree. Moreover, some techniques focus on sampling subgraphs in each mini-batch, like GraphSAINT [47] and ClusterGCN [10]. While these techniques effectively scale GNNs to larger graphs, they do not adapt to the sampling policy based on the GNN’s performance on the task. In graph signal processing community, the authors in [15] propose a node sampling technique based on [4] that aims to sample nodes for a unique and stable graph signal reconstruction. However, this method is only applied to small graphs.
@@
Learnable Sampling Policy: A few methods learn the probability of including a node based on feedback from the GNN. AS-GCN [18] is a method that learns a linear function that estimates the node probabilities layer-wise. Similarly, PASS [44] learns a mixture of a random distribution and a learned policy with RL. Like GRAPES, it is task-adaptive, however, its sampling probability uses a bilinear similarity measure between neighboring node features, meaning the graph structure is not considered. GNN-BS [27] formulates the node-wise sampling problem as a bandit problem and updates the sampling policy according to a reward function that reduces the sampling variance. SubMix [1] proposes a mixture distribution of sampling heuristics with learnable mixture weights. DSKReG [40] learns the relevance of items in a user-item knowledge graph by jointly optimizing the sampling strategy and the recommender model. The majority of these methods focus on variance reduction and fail to consider the classification loss, unlike GRAPES. We argue that adaptivity to the classification loss, allows for sampling the influential nodes depending on the task, and results in better performance.
@@
Other Scalable Methods: Authors of [33] propose a method to transfer the weights of a GNN trained on a mid-sized graph to larger graphs given the graphon similarity between the graphs. Another group of papers uses historical embeddings of the nodes when updating the target nodes’ embeddings [8, 13, 45, 37]. GAS [13] approximates the embeddings of the 1-hop neighbors using the historical embeddings of those nodes learned in the previous training iterations. These methods reduce the GPU memory usage by training in mini-batches and learning from the 1-hops neighbors with the historical embeddings saved in CPU memory. Unlike GRAPES, they process the whole graph.
@@ 3 Background: GNN Training and Sampling
We first provide the necessary background about GNNs and graph sampling. Although our method is independent of the choice of GNN architecture, we limit our discussion to the GCN architecture for simplicity [23].
Let $\mathcal{G}=(\mathcal{V,E})$ be an undirected graph with a list of $N$ nodes $\mathcal{V}=\{1,\ldots,N\}$ and a set of edges $\mathcal{E}$. The adjacency matrix $A\in\{0,1\}^{N\times N}$ indicates a connection between a pair of nodes. Next, let $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ and where $\tilde{D}$ is the degree matrix of $\tilde{A}$. Let $X\in\mathbb{R}^{N\times f}$ be the node embeddings and let $Y$ be the labels for the target nodes $\mathcal{V}^{t}\subset\mathcal{V}$, where $\mathcal{V}^{t}$ indexes the nodes with a label. In node classification, we want to learn to predict the labels of the target nodes.
@@
As the number of layers increases, the computation of the embedding of the node $i$ involves neighbors from further hops. As a result, the neighborhood size grows rapidly with the number of layers. We study how to sample the graph to overcome this growth. We focus on layer-wise sampling, which is a common type of graph sampling approaches [9, 51, 18]. First, we divide the target nodes in mini-batches of size $b$. Then, in each layer, we sample $k$ nodes $\mathcal{V}^{(l)}$ among the neighbors of the nodes in the previous layer using the sampling policy $q$. To make this precise, we will need some additional notation. The computation of the output of layer $l$ is then:
@@
    $K^{(0)}=\mathcal{V}^{(0)}$ is the set of target nodes in the current mini-batch (a subset of $\mathcal{V}^{t}$);
@@
Existing layer-wise sampling methods such as LADIES or FastGCN use a fixed heuristic to determine $q$, for instance, by computing node probabilities proportional to the node degrees. However, an adaptive method learns the distribution $q$ instead.
@@ 4 Graph Adaptive Neighbor Sampling (GRAPES)
In this section, we introduce Graph Adaptive Sampling (GRAPES). GRAPES is a layer-wise and layer-dependent sampling method. In each layer $l$, we sample a subset $\mathcal{V}^{(l)}$ that is much smaller than the neighborhood of the previous layer. That is, $|\mathcal{V}^{(l)}|=k\ll|\mathcal{N}(K^{(l-1)})|$, where again $K^{(l-1)}=\mathcal{V}^{(0)}\cup\mathcal{V}^{(l-1)}$ adds the batch nodes to the sampled nodes of the previous layer. We use a second GNN to compute the inclusion probability for each node in $\mathcal{N}(K^{(l-1)})$. In the remainder of this section, we describe our sampling policy $q$ and training methods for the sampling policy. Furthermore, Algorithm 1 shows one epoch of GRAPES in pseudocode.
@@ 4.2 Sampling exactly $k$ nodes
Instead, we use the Gumbel-Top-$k$ trick [39, 19], which selects a set of exactly $k$ nodes $\mathcal{V}^{(l)}$ by perturbing the log probabilities randomly and taking the top-$k$ among those:
@@
While this results in a tractable and adaptive sampling procedure of exactly $k$ nodes for the classifier GNN, we have not yet given a method for learning the sampling policy. Unfortunately, a sampling operation, or in this case, the top-$k$ operation, provides no functional gradient [30]. We will resort to simple methods from the reinforcement learning and GFlowNet literature to still be able to train the sampling policy, which we will explain next.
@@ 4.3 Training the Sampling Policy
We train the sampling policy to minimize the classification loss $\mathcal{L}_{\operatorname{C}}$ of the classifier GNN [^2] GRAPES can be extended to other tasks, but we focus on node classification in the current work.. We use two methods to train the sampling policy $\operatorname{GCN}_{\operatorname{S}}$: a reinforcement learning (RL) method and a GFlowNet (GFN) method.
@@
In the REINFORCE-based method, we use a simple REINFORCE estimator [41] to compute an unbiased gradient of the classification loss.
@@
where we sample $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}\sim q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)},k)$. Taking the derivative with respect to the parameters of $q$ results in the standard REINFORCE estimator. Note that this is an off-policy estimator since we sample from $q$ conditioned on the number of samples $k$ with the Gumbel-Top-$k$ trick, but compute gradients with respect to the distribution unconditioned on $k$. This is because computing likelihoods conditioned on $k$, although possible [3], is computationally expensive. We discuss this issue in more detail in Appendix A.
@@
The second method uses the Trajectory Balance loss [28] from the GFlowNet literature [7, 6], which is known to perform well in off-policy settings [29].

### v3

@@ Abstract
Graph neural networks (GNNs) learn to represent nodes by aggregating information from their neighbors. As GNNs increase in depth, their receptive field grows exponentially, leading to high memory costs. Several works in the literature proposed to address this shortcoming by sampling subgraphs or by using historical embeddings. These methods have mostly focused on benchmarks of single-label node classification on homophilous graphs, where neighboring nodes often share the same label. However, most of these methods rely on static heuristics that may not generalize across different graphs or tasks. We argue that the sampling method should be adaptive, adjusting to the complex structural properties of each graph. To this end, we introduce GRAPES, an adaptive sampling method that learns to identify the set of nodes crucial for training a GNN. GRAPES trains a second GNN to predict node sampling probabilities by optimizing the downstream task objective. We evaluate GRAPES on various node classification benchmarks involving homophilous as well as heterophilous graphs. We demonstrate GRAPES’ effectiveness in accuracy and scalability, particularly in multi-label heterophilous graphs. Additionally, GRAPES uses orders of magnitude less GPU memory than a strong baseline based on historical embeddings. Unlike other sampling methods, GRAPES maintains high accuracy even with smaller sample sizes and, therefore, can scale to massive graphs. Our implementation is publicly available online. [^1] Available at https://github.com/dfdazac/grapes.
@@ 1 Introduction
Despite the broad range of applications of GNNs (Nettleton, 2013; Wu et al., 2022; Li et al., 2022; Kipf & Welling, 2016; Velickovic et al., 2017; Yun et al., 2019), scalability remains a significant challenge (Serafini & Guan, 2021). Unlike traditional machine learning problems, where data is assumed to be i.i.d., the graph structure introduces dependencies between a node and its neighborhood. This complicates partitioning data into mini-batches. Additionally, the number of nodes that a GNN needs to process increases exponentially with the number of layers (Ma et al., 2025).
Most current graph sampling methods use a fixed heuristic to compute node inclusion probabilities: The sampling process is independent of the node features, graph structure, or the GNN task performance (Chen et al., 2018b; Zou et al., 2019; Zeng et al., 2019). For simple graphs like homophilous graphs (Zhu et al., 2020; Zheng et al., 2022; Zhao et al., 2023) where there is a strong correlation between labels of a node and its neighbors, a fixed heuristic typically suffices. However, these methods fall short when applied to graphs in which only a few of the nodes in the neighborhood provide relevant information for the task at hand. Consequently, we study adaptive sampling, where the sampling method adapts dynamically to the task by learning which nodes should be included. Our adaptive sampling technique learns to sample by directly minimizing the classification loss of the downstream GNN. This differs from most existing graph sampling methods, which aim to approximate the full-batch GNN (Chen et al., 2018b; Zou et al., 2019; Zeng et al., 2019; Huang et al., 2018), opting for an indirect approach to achieve high accuracy.
@@
To enable backpropagation through the sampling step, we need a gradient estimator. We compare a reinforcement learning (RL) approach and a GFlowNet (GFN) approach (Bengio et al., 2021a). Both approaches allow us to consider the sampling process and the GNN computation together and to train them concurrently. This allows the sampler to adapt based on factors like the node features, the graph structure, the sample size, and other contextual features.
In multi-class classification– where a node can be assigned only one of multiple classes– on homophilous graphs, even a simple strategy like random sampling could be sufficient, since the neighborhood of a node contains nodes with similar labels (see Fig. 2(a)). In the case of multi-label classification, on the other hand, the number of possible labels assigned to a node is $|2^{\mathcal{Y}}|$, where $\mathcal{Y}$ is the unique set of labels. Furthermore, in heterophilous graphs, the relationship between the neighborhood of a node and its label is more complex (see Fig. 2(b)). We argue that in such cases, a sampling policy that directly adapts to the properties of the graph and the downstream task is crucial.
Therefore, in addition to the datasets commonly used in the literature, we evaluate GRAPES on several heterophilous and/or multi-label graphs and demonstrate its effectiveness on these complicated graphs. Moreover, we spend significant effort on the evaluation protocol for fairly testing graph sampling methods. In particular, we ensure that all sampling methods are evaluated under the same conditions and on the same GNN architecture to eliminate any confounding factors and to ensure that any changes in performance can only be attributed to the sampling method. To the best of our knowledge, we are the first to perform such a rigorous comparison on twelve varied datasets.
Additionally, we provide a theoretical analysis indicating the effectiveness of adaptive sampling on a specific category of heterophilous graphs, where only a subset of neighbors contains relevant information for an accurate node classification.
@@
    Performance. GRAPES achieves state-of-the-art performance in multi-label classification benchmarks over heterophilous graphs, while performing competitively on multi-class classification on homophilous graphs.
@@
    Memory Efficiency. GRAPES uses orders of magnitude less GPU memory when compared with a strong baseline that relies on historical embeddings for scaling GNN training.
@@
    Robustness. In comparison with other methods, GRAPES maintains good performance under exponential reductions in sample size.
@@ 2 Related Work
Fixed Sampling Policy: In this category, the sampling policy is independent of the training of the GNN and is based on a fixed heuristic that does not involve any training. Given a set of target nodes, i.e. the nodes that are to be classified, node-wise sampling methods sample a given number of nodes for each target node. GraphSage (Hamilton et al., 2017) is a node-wise sampling method that randomly samples nodes. However, node-wise sampling can result in nodes being sampled multiple times redundantly because they can be the neighbors of several nodes (Zou et al., 2019). A more efficient approach is layer-wise sampling, for example, FastGCN (Chen et al., 2018b) and LADIES (Zou et al., 2019). They aim to minimize variance by sampling nodes in each layer with probabilities proportional to their degree. MVS-GNN (Cong et al., 2020) proposes a fixed sampling policy with the aim of minimizing sampling variance by decoupling it into two components. The first component, embedding approximation variance, arises from neighbor sampling and is mitigated through the use of historical embeddings. The second component, stochastic gradient variance, results from mini-batching and is addressed by incorporating the norm of the node gradients into the sampling process. Moreover, some techniques focus on sampling subgraphs in each mini-batch, like GraphSAINT (Zeng et al., 2019) and ClusterGCN (Chiang et al., 2019). While these techniques effectively scale GNNs to larger graphs, they do not adapt to the sampling policy based on the GNN’s performance on the task. In the graph signal processing community, the authors in (Geng et al., 2023) propose a node sampling technique based on (Anis et al., 2016) to sample nodes for a unique and stable graph signal reconstruction. However, this method is only applied to small graphs.
Figure 2: The downstream task and the structural properties of the graph affect the strategy used for sampling. Fig. 2(a) shows a citation network where articles citing each other are likely to belong to the same category. Fig. 2(b) shows a graph of interacting proteins, each with different biological functions.
@@
Learnable Sampling Policy: A few methods learn the probability of including a node based on feedback from the GNN. AS-GCN (Huang et al., 2018) is a method that learns a linear function that estimates the node probabilities layer-wise. Similarly, PASS (Yoon et al., 2021) learns a mixture of a random distribution and a learned policy with RL. Like GRAPES, it is task-adaptive; however, its sampling probability uses a bilinear similarity measure between neighboring node features, meaning the graph structure is not considered. FairSample Cong et al. (2023) is a learnable sampling policy that borrows the PASS training mechanism that utilizes reinforcement learning to combine classification loss and a fairness loss to balance accuracy and fairness towards the underrepresented nodes. GNN-BS (Liu et al., 2020) formulates the node-wise sampling problem as a bandit problem and updates the sampling policy according to a reward function that reduces the sampling variance. SubMix (Abu-El-Haija et al., 2023) proposes a mixture distribution of sampling heuristics with learnable mixture weights. DSKReG (Wang et al., 2021) learns the relevance of items in a user-item knowledge graph by jointly optimizing the sampling strategy and the recommender model. The majority of these methods focus on variance reduction and fail to consider the classification loss, unlike GRAPES. We argue that adaptivity to the classification loss allows for sampling the influential nodes depending on the task and results in better performance.
@@
Other Scalable Methods: Authors of (You et al., 2022; Zhang et al., 2024) use graph lottery tickets to eliminate unnecessary edges, while DSpar (Liu et al., 2023) uses a simple degree-based heuristic to sparsify the graph. Ruiz et al. (2023) proposes a method to transfer the weights of a GNN trained on a mid-sized graph to larger graphs, given the graphon similarity between the graphs. Another group of papers uses historical embeddings of the nodes when updating the target nodes’ embeddings (Chen et al., 2018a; Fey et al., 2021; Yu et al., 2022; Shi et al., 2023). GAS (Fey et al., 2021) approximates the embeddings of the 1-hop neighbors using the historical embeddings of those nodes learned in the previous training iterations. These methods reduce the GPU memory usage by training in mini-batches and learning from the 1-hop neighbors with the historical embeddings saved in CPU memory. Unlike GRAPES, they process the whole graph.
@@ 3 Background: GNN Training and Sampling
We first provide the necessary background about GNNs and graph sampling. Although our method is independent of the choice of GNN architecture, we limit our discussion to the GCN architecture for simplicity (Kipf & Welling, 2016).
Let $\mathcal{G}=(\mathcal{V,E})$ be an undirected graph with a list of $N$ nodes $\mathcal{V}=\{1,\ldots,N\}$ and a set of edges $\mathcal{E}$. The adjacency matrix $A\in\{0,1\}^{N\times N}$ indicates a connection between a pair of nodes. Next, let $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ and where $\tilde{D}$ is the degree matrix of $\tilde{A}$. Let $X\in\mathbb{R}^{N\times f}$ be the node embeddings and let $Y$ be the labels for the target nodes $\mathcal{V}^{t}\subset\mathcal{V}$, where $\mathcal{V}^{t}$ indexes the nodes with a label.
@@
As the number of layers increases, the computation of the embedding of the node $i$ involves neighbors from further hops. As a result, the neighborhood size grows rapidly with the number of layers. We study how to sample the graph to overcome this growth. We focus on layer-wise sampling, a common type of graph sampling approach (Chen et al., 2018b; Zou et al., 2019; Huang et al., 2018). First, we divide the target nodes into mini-batches of size $b$. Then, in each layer, we sample $k$ nodes $\mathcal{V}^{(l)}$ among the neighbors of the nodes in the previous layer using the sampling policy $q$. To make this precise, we will need some additional notation. The computation of the output of layer $l$ is then:
@@
    $K^{(0)}=\mathcal{V}^{(0)}$ is the set of target nodes in the current mini-batch, where $\mathcal{V}^{(0)}\subset\mathcal{V}^{t}$;
@@
Existing layer-wise sampling methods, such as LADIES or FastGCN, use a fixed heuristic to determine $q$, for instance, by computing node probabilities proportional to the node degrees. However, an adaptive method learns the distribution $q$ instead.
@@ 4 Graph Adaptive Neighbor Sampling (GRAPES)
We introduce Graph Adaptive Sampling (GRAPES). GRAPES is a layer-wise and layer-dependent adaptive sampling method that learns a sampling policy that minimizes the training objective conditioned on the input graph. In each layer $l$, we sample a subset $\mathcal{V}^{(l)}$ that is much smaller than the neighborhood of the previous layer. That is, $|\mathcal{V}^{(l)}|=k\ll|\mathcal{N}(K^{(l-1)})|$, where again $K^{(l-1)}=\mathcal{V}^{(0)}\cup\mathcal{V}^{(l-1)}$ adds the batch nodes to the sampled nodes of the previous layer. We use a second GNN to compute the inclusion probability for each node in $\mathcal{N}(K^{(l-1)})$. In the remainder of this section, we describe our sampling policy $q$ and training methods for the sampling policy. Furthermore, Algorithm 1 shows one epoch of GRAPES in pseudocode.
@@ 4.2 Sampling exactly k nodes
Instead, we use the Gumbel-Top-$k$ trick (Vieira, 2014; Huijben et al., 2022), which selects a set of exactly $k$ nodes $\mathcal{V}^{(l)}$ by perturbing the log probabilities randomly and taking the top-$k$ among those:
@@
While this results in a tractable and adaptive sampling procedure of exactly $k$ nodes for the classifier GNN, we have not yet given a method for learning the sampling policy. Unfortunately, a sampling operation, or in this case, the top-$k$ operation, provides no functional gradient (Mohamed et al., 2020). We will resort to simple methods from the reinforcement learning and GFlowNet literature to still be able to train the sampling policy, which we will explain next.
The choice of a Bernoulli distribution for sampling node neighborhoods is motivated by the idea that when conditioned on a target node, we can sample a node independent of other nodes in the neighborhood to maintain efficiency. Other assumptions could be more elaborate, such as sampling jointly pairs of nodes at a time, but this would in turn increase the computational cost of computing probabilities and sampling. On the other hand, the Gumbel top-k trick allows us to sample exactly $k$ elements without replacement, and while there are other applicable methods such as reservoir sampling, they are mathematically equivalent (Huijben et al., 2022).
@@ 4.3 Training the Sampling Policy
We train the sampling policy to minimize the classification loss $\mathcal{L}_{\operatorname{C}}$ of the classifier GNN [^3] GRAPES can be extended to other tasks, but we focus on node classification in the current work.. We use two methods to train the sampling policy $\operatorname{GCN}_{\operatorname{S}}$: a reinforcement learning (RL) method and a GFlowNet (GFN) method. We also experimented with the straight-through estimator (Bengio et al., 2013) for learning in GRAPES, which we observed not to perform well due to increased computation and high bias in the estimated gradients (see Appendix G).
@@
In the REINFORCE-based method, we use a simple REINFORCE estimator (Williams, 1992) to compute an unbiased gradient of the classification loss.
@@
where we sample $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}\sim q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)},k)$. Taking the derivative with respect to the parameters of $q$ results in the standard REINFORCE estimator. Note that this is an off-policy estimator since we sample from $q$ conditioned on the number of samples $k$ with the Gumbel-Top-$k$ trick, but compute gradients with respect to the distribution unconditioned on $k$. This is because computing likelihoods conditioned on $k$, although possible (Ahmed et al., 2023), is computationally expensive. We discuss this issue in more detail in Appendix A.
@@
The second method uses the Trajectory Balance loss (Malkin et al., 2022a) from the GFlowNet literature (Bengio et al., 2021b; a), which is known to perform well in off-policy settings (Malkin et al., 2022b).
@@
Memory complexity.
Layer-wise sampling methods like GRAPES, FastGCN, and LADIES have a sampling space complexity of $O(Dbk)$, where $D$ is the maximum node degree, $b$ is the batch size, $k$ is the sample size, and $L$ is the number of layers. In GAS time complexity is $O(bk)$ due to the use of historical embeddings, and memory complexity is $O(DbL+N)$ where $N$ is the extra overhead for storing historical embeddings for the $N$ nodes in the graph, which can be significantly larger than $DbL$.
@@
5 Theoretical Analysis
In this section, we provide theoretical insights on the performance difference between adaptive and non-adaptive samplers, and we prove that adaptive sampling can achieve higher accuracy than non-adaptive sampling for a specific category of heterophilous graphs. We show that if there exists specific crucial information among the neighbors, there are adaptive sampling methods that perfectly identify these neighbors, while non-adaptive sampling methods cannot distinguish them from the other neighbors.
Let $\mathcal{G}=(\mathcal{V,E)}$ be an undirected graph with a set of N nodes $\mathcal{V}=\{1,\dots,N\}$ and a set of edges =$\mathcal{E}=\{e_{ij}\}_{i,j=1}^{N}$ where $e_{ij}\text{ denotes an edge between }v_{i}\text{ and }v_{j}$. Let $X\in\mathbb{R}^{N\times F}$ be the node features and let $Y=\{y_{i}\}_{i=1}^{N}$ be a set of node labels. The adjacency matrix $A\in\{0,1\}^{N\times N}$ indicates a connection between a pair of nodes. Let $\mathcal{N}(v_{i})=\{v_{j}|e_{ij}\in\mathcal{E}\}$ indicate the set of $v_{i}$’s neighbors and $\mathcal{N}^{L}(v_{i})=\{v_{j}|l(v_{i},v_{j})\leq L\}$ indicate the $L$-hop neighborhood of $v_{i}$, where $l(v_{i},v_{j})$ indicate the shortest path between $v_{i}$ and $v_{j}$. In the following definitions and theorems, for simplicity, we consider the batch size of one.
Definition 1 (Neigbor sampling).
A neighbor sampling method is a method that assigns scores to neighbors of a target node, and for a given sampling budget $K$, samples $K$ neighbors proportional to their score. An $L$-layer neighbor sampling method, for each layer $l$, assigns scores to the l-hop neighbors of a target node.
Definition 2 (Featureless and feature-based sampling).
An $L$-layer feature-less sampling method is a neighbor sampling method in which the sampling score for each node $v_{i}$ is a function of only its $L$-hop neighborhood, i.e. $s(v_{i})=g(v_{i},\mathcal{N}^{L}(v_{i}))$, where $s(v_{i})$ is the sampling for node $v_{i}$ and $g$ is the sampling function. In an $L$-layer feature-based sampling method, the sampling function is a function of the node features in addition to the $L$-hop neighborhood, that is, $s(v_{i})=g^{\prime}(v_{i},X,\mathcal{N}^{L}(v_{i}))$.
In practice, all feature-based sampling methods are adaptive since the sampling methods learns the sampling score from the features, and all non-adaptive methods are featureless. Therefore, in this section, we refer to the sampling methods as adaptive and non-adaptive.
Random sampling is a non-adaptive sampler with a constant sampling score. LADIES and FastGCN are non-adaptive samplers whose sampling scores are proportional to the node degrees. GraphSAINT-RW, i.e., the random walk-based sampler of GraphSAINT, which we used in our experiments, is a non-adaptive sampler where the score correlates with how frequently a node appears in $L$-hop neighborhoods, influenced by both the node’s degree and its position in the graph. AS-GCN, PASS, and GRAPES are adaptive samplers.
Definition 3 (Zhu et al. (2020)).
Edge homophily is the ratio of edges that connect two nodes of the same label:
$h=\frac{|\{e_{ij}\in\mathcal{E}:y_{i}=y_{j}\}|}{|\mathcal{E}|}$
In the next theorem, we show that the optimal adaptive and non-adaptive samplers perform differently for a certain category of graphs, with adaptive samplers perfectly sampling the influential neighbors and non-adaptive samplers sampling such nodes with a vanishing probability.
Theorem 1.
There exist undirected graphs such that a GCN with an $L$-layer adaptive sampler and a sampling rate of $K$ neighbors per node per layer can perform with perfect accuracy, while GCNs with a non-adaptive sampler can achieve an accuracy higher than chance level only with probability $p\leq\frac{LK}{N}$.
For the proof of this theorem, please refer to Appendix H. In the proof, we construct a family of distributions over graphs where the presence of a single neighbor is essential for a correct classification of a target node. Then, we show that only feature-based samplers, by learning from the node features, can sample those neighbors, while featureless samplers, with overwhelming probability, cannot. Since, in practice, all feature-based samplers are adaptive and, inversely, all non-adaptive samplers are featureless, we conclude that adaptive models can solve this task, while non-adaptive methods cannot.
Additionally, we show that such graphs are heterophilous. Moreover, often, multi-label graphs are heterophilous because neighboring nodes are less likely to share the exact same set of labels. Even slight variations in the labels of neighboring nodes can lead to heterophily. Therefore, adaptive sampling methods outperform non-adaptive ones in such graphs.